One mole of an ideal gas expands isothermally to five times its initial volume.
Calculate the entropy change in terms of R, the gas constant.
1
Expert's answer
2017-03-24T13:54:07-0400
S = nRln(V2/V1) = 1 mol * R * ln(5) = 1.61R J*mol-1K-1